Determine the concentration of hydroxide ions for a 25∘C solution with a pOH of 4.56.

  • Enter your answer with 2 significant figures.

Sorry, that's incorrect. Try again?

3.6x10^-10
Answer #1

Given that;

pOH = 4.56

We know that;

pOH = - log [OH-]

[OH-] = 10^(-pOH) = 10^(-4.56) = 2.75*10^-5 M  

OR

[OH-] = 2.8*10^-5 M .....Answer

Let me know if any doubts.